# Graph Knowledge Distillation to Mixture of Experts

### v1

@@ 4.3 MoE initialization
In most implementations, the embeddings of the experts are initialised randomly (Chi et al., 2022; Li et al., 2022; Yan & Li, 2023). In our approach, we desire more informative initialisations, because embeddings are operating in the input space. We therefore apply a pretraining stage, following Zhang et al. (2021a). We pretrain the model for several epochs, routing all input vectors to the first expert. Subsequently, we clone the parameters of the first expert to all other experts. We reset the optimiser state after pretraining. We collect the inputs for each MoE layer and apply $L_{2}$ normalization. We then apply K-means clustering, with the number of clusters equal to the number of experts. We initialise the embeddings with the cluster centers.
@@ 5.1 Experimental setting
Model. Our Routing by Memory models have the same number of layers as the teacher. Every expert is a linear layer with the same size as the corresponding layer of the teacher. We use up to 8 experts for RbM (the exact number for each dataset is selected using the validation set). We use the same number of experts for all the RbM layers inside the model. Three experts are active at a time. RbM routing is sensitive to dropout, so we avoid the application of dropout directly before the RbM layers. In our model, we apply dropout before the input of the expert for each layer except the first one.
@@ 5.4 Ablation study, label propagation, and number of experts
Table 5: Ablation study on equation 11 loss components. GraphSAGE teacher is used. Results are showing accuracy (higher is better). RbM performance is provided for the reference. For “KD only” column we are reporting results with KD and KDR losses only and no embedding losses.
@@
Loss terms. We now examine whether each component of the equation 11 is important for achieving better performance. Our model includes three additional loss terms for each RbM layer (see equation 11): commitment loss (equation 6), self-similarity loss (equation 7), and load balance loss (equation 8). In order to conduct the ablation study we remove each component individually. None of the other hyperparameters is changed. The results are presented in Table 5, and show that removing any of the loss components reduces performance. The performance deterioration is relatively small for each dataset, but it is observed for every setting and every loss. It is not clear that any of the three loss terms is the most important. Any combination that includes two loss terms outperforms (by a small margin) the baseline with all three removed for all medium and large datasets, indicating that all three loss terms contribute.
@@ 5.5 Routing spatial structure analysis
(a) Plain MoE internal representation colored by top expert (left) and by label (right).
(b) RbM internal representation colored by top expert (left) and by label (right).
@@
In order to analyse the routing spatial structure qualitatively, we utilise the T-SNE (Van der Maaten & Hinton, 2008), with perplexity of 30 and PCA initialization, to produce a 2-d visualizations of a router embedding space for RbM and a vanilla MoE in Figure 4. These correspond to the hidden representation $h$ for RbM (see Equation 3) and the linear projection of the hidden representation, $Wh$, for the MoE ( see Equation 2). We trained both MoE and RbM models on the Academic-Physics dataset in the transductive setting with the same teacher and selected the router of the last layer to produce the representation. In Figure 4, the hidden representations are colored according to the labels (on the left) and according to the top-score expert (on the right). Figure 4 shows that MoE experts mix and distribute datapoints of the same classes between different experts, while RbM experts have a clear specialization.
@@ 6 Conclusion and Future work
In this paper we focused on the task of distillation from a graph neural network and introduced RbM, a Mixture of Experts model that encourages strong expert specialization at the routing level. We established how parameter inflation can positively affect the performance and showed practical application of MoE in the knowledge distillation domain. Our approach outperforms existing baselines on most medium-size or large datasets. The key innovations of our approach is in embeddings to be the part of the hidden representation space.
We used additional losses hidden space is forced into the shape of multiple elliptical clusters, which could be too restrictive and thus suboptimal. In order to reduce number of hyperparameres we assumed all layer of RbM to have the same number of experts and thus clustered into the same number of clusters. Selecting a suitable number of experts for the layer automatically can improve the performance of the overall model. Both off these assumptions led to the model being more sensitive to the number of experts than the plain MoE. We leave those direction of improvement as a future work. In addition to that a a future direction can be about the application of the MoE in graph domain. This work uses a concatenation of positional and feature vector for both routing and selected experts processing, however an alternative approach can be to route with positional vector while feature vector is supplied to the selected expert.
@@ Appendix B Hardware specification
Our experiments were conducted using an NVIDIA Tesla V100 GPU with 32GB of memory. The machine has an Intel Xeon Gold 6140 CPU with clock frequency of 2.30GHz and total thread count of 36. All computations, with exception of the clustering, were executed on the GPU. For Cora, Citeseer, PubMed, Amazon-Comp, Amazon-Photo and Academic-CS datasets we executed five parallel runs simultaneously. Each parallel run was allocated 6GB of GPU memory and 5 threads for the clustering. For Academic-Phy, OGB-ArXive, OGB-Products we executed only one run at a time with 32GB of GPU memory and 10 threads for clustering. We were unable to run KRD+ on our setup as it requires more than 32GB of memory.
Appendix C Hyperparameters tuning protocol
We are using Ray Tune (Liaw et al., 2018) to tune model hyperparameters. We tuned the following model structure hyperparameters: (i) dropout rate was selected from $[0.0,0.1,0.2,0.3,0.4,0.5,0.6]$ and applied to all dropout layers in the model; (ii) total number of experts was selected $[4,5,6,7,8]$. In addition to the structure hyperparameters we selected the following training hyperparameters: (i) learning rate for Adam optimizer (Kingma & Ba, 2014) was chosen from $[0.01,0.005,0.001]$; (ii) weight $\alpha$ of the commitment loss (6) from the range $[0.0,0.1]$; (iii) weights $\beta$ and $\gamma$ of the the load-balancing loss (8) and self-similarity loss 7 correspondingly from the range $[0.0,0.05]$.
@@ Appendix F Complexity analysis
Table 9: Complexity analysis for the student models. Since we are using sparse routing, we present active parameter counter along with the total parameters. Only active parameters are participating in the inference. Note that these empirical results are specific to our setup and are not generalised.

### v2

@@ 4.3 MoE initialization
In most implementations, the embeddings of the experts are initialised randomly (Chi et al., 2022; Li et al., 2022; Yan & Li, 2023). In our approach, we desire more informative initialisation, because embeddings are operating in the input space. We therefore apply a pretraining stage, following Zhang et al. (2021a). We pretrain the model for several epochs, routing all input vectors to the first expert. Subsequently, we clone the parameters of the first expert to all other experts. We reset the optimiser state after pretraining. We collect the inputs for each MoE layer and apply $L_{2}$ normalization. We then apply K-means clustering, with the number of clusters equal to the number of experts. We initialise the embeddings with the cluster centers.
Operating directly in the space of the hidden representations not only removes the superfluous parameters of the projection matrix from the model, but it also makes initialization of the experts much easier. This helps to avoid representation collapse that can be caused by randomly initialized projection matrix.
@@ 5.1 Experimental setting
Model. Our Routing-by-Memory models have the same number of layers as the teacher. Every expert is a linear layer with the same size as the corresponding layer of the teacher. We use up to 8 experts for RbM (the exact number for each dataset is selected using the validation set). We use the same number of experts for all the RbM layers inside the model. Three experts are active at a time. RbM routing is sensitive to dropout, so we avoid the application of dropout directly before the RbM layers. In our model, we apply dropout before the input of the expert for each layer except the first one. We use a linear annealing schedule for the updating of the expert embeddings (see Appendix G).
@@ 5.2 Performance comparison
Since we perform pretraining and clustering in our model initialization, there is some additional computational overhead compared to training parameter-inflated baselines. In our experiments, pretraining does not exceed 7% of the total number of training epochs on average (in our implementation we use early stop, so the number of epochs can vary). The clustering overhead is effectively insignificant. On our setup, it takes 80 seconds to cluster all required embeddings for RbM on OGB-ArXive, one of the larger datasets.
@@ 5.4 Ablation study, label propagation, and number of experts
Table 5: Ablation study on equation 13 loss components. GraphSAGE teacher is used. Results are showing accuracy (higher is better). RbM performance is provided for reference. For “KD only” column we report results with KD and KDR losses only and no embedding losses.
@@
Loss terms. We now examine whether each component of the equation 13 is important for achieving better performance. Our model includes three additional loss terms for each RbM layer (see equation 13): commitment loss (6), self-similarity loss (7), and load balance loss (8). In order to conduct the ablation study we remove each component individually. None of the other hyperparameters is changed. The results are presented in Table 5, and show that removing any of the loss components reduces performance. The performance deterioration is relatively small for each dataset, but it is observed for every setting and every loss. It is not clear that any of the three loss terms is the most important. Any combination that includes two loss terms outperforms (by a small margin) the baseline with all three removed for all medium and large datasets, indicating that all three loss terms contribute.
@@ 5.5 Routing spatial structure analysis
((a)) Plain MoE internal representation colored by top expert (left) and by label (right).
((b)) RbM internal representation colored by top expert (left) and by label (right).
@@
In order to analyse the routing spatial structure qualitatively, we utilise the T-SNE (Van der Maaten & Hinton, 2008), with perplexity of 30 and PCA initialization, to produce 2-d visualizations of the router embedding space for RbM and a vanilla MoE in Figure 4. These correspond to the hidden representation $h$ for RbM (see equation 3) and the linear projection of the hidden representation, $Wh$, for the MoE (see equation 2). We trained both MoE and RbM models on the Academic-Physics dataset in the transductive setting with the same teacher and selected the router of the last layer to produce the representation. In Figure 4, the hidden representations are colored according to the labels (on the left) and according to the top-score expert (on the right). Figure 4 shows that MoE experts mix and distribute datapoints of the same classes between different experts, while RbM experts have a clear specialization.
@@ 6 Conclusion and Future work
In this paper we focused on the task of distillation from a graph neural network and introduced RbM, a Mixture of Experts model that encourages strong expert specialization at the routing level. We established how parameter inflation can positively affect the performance and showed practical application of MoE in the knowledge distillation domain. Our approach outperforms existing baselines on most medium-size or large datasets. A key innovation in our approach is the choice to have expert embeddings reside in the hidden representation space. We incorporate additional loss terms to improve specialization and load balancing.
One limitation of our work is the use of K-means clustering, which implies that we identify elliptical clusters during expert initialization. This could be overly restrictive and thus suboptimal. Also, in order to reduce the number of hyperparameters, we use the same number of experts for every layer. Selecting a suitable number of experts for each layer automatically could potentially improve the overall performance. Due to these choices, our proposed model is more sensitive to the number of experts than vanilla MoE, although a suitable number of experts can be identified using a small validation set. Our approach uses a concatenation of positional encoding and feature vector for both routing and processing by the selected experts. An alternative approach deserving investigation is to route using the positional encoding and providing only the feature vector to the selected expert.
@@ Appendix B Hardware specification
Our experiments were conducted using an NVIDIA Tesla V100 GPU with 32GB of memory. The machine has an Intel Xeon Gold 6140 CPU with clock frequency of 2.30GHz and total thread count of 36. All computations, with the exception of the clustering, were executed on the GPU. For Cora, Citeseer, PubMed, Amazon-Comp, Amazon-Photo and Academic-CS datasets we executed five parallel runs simultaneously. Each parallel run was allocated 6GB of GPU memory and 5 threads for the clustering. For Academic-Phy, OGB-ArXive, OGB-Products we executed only one run at a time with 32GB of GPU memory and 10 threads for clustering. We were unable to run KRD+ on our setup as it requires more than 32GB of memory.
Appendix C Hyperparameter tuning
We used Ray Tune (Liaw et al., 2018) to tune model hyperparameters. Specifically, we used the Optuna search algorithm (Akiba et al., 2019). We sampled 200 hyperparameter configurations for the small and medium datasets and 80 for the large datasets. We tuned the following model structure hyperparameters: (i) dropout rate was selected from $[0.0,0.1,0.2,0.3,0.4,0.5,0.6]$ and applied to all dropout layers in the model; (ii) total number of experts was selected from $[4,5,6,7,8]$. In addition to the structure hyperparameters, we selected the following training hyperparameters: (i) learning rate for the Adam optimizer (Kingma & Ba, 2014) was chosen from $[0.01,0.005,0.001]$; (ii) weight $\alpha$ of the commitment loss (6) from the range $[0.0,0.1]$; (iii) weight $\beta$ of the load-balancing loss (8) and weight $\gamma$ of the self-similarity loss (7) both from the range $[0.0,0.05]$.
Compared to a distillation setup with an MLP student, our training setup introduces three loss hyperparameters (see equation (13)) and an embedding update hyperparameter (see equation (4)). Due to the usage of hyperparameter sampling, these additional hyperparameters do not noticeably increase the search time.
@@ Appendix F Complexity analysis
Table 9: Complexity analysis for the student models. Since we are using sparse routing, we present active parameter counter along with the total parameters. Only active parameters are participating in the inference.
@@
Appendix G Annealing schedule
In order to ensure the convergence of expert embeddings, we apply a linear annealing schedule:
$\lambda(t)=\lambda_{0}+\frac{(1-\lambda_{0})\Delta}{T}t,$ (14)
where $10\leq\Delta<$ 1 is an annealing hyperparameter, $T$ is the expected number of epochs and $\lambda_{0}$ is a constant that specifies the initial update rate. In our experiments, we set $\lambda_{0}=0.9$, $T=200$ and $\Delta=0.05$.
With the following experiment, we demonstrate how varying $\Delta$ affects the performance of RbM. We repeat the inductive and transductive experiments for the OGB-ArXive dataset using GraphSAGE as a teacher model, but vary the annealing parameter $\Delta$. From Table 10, we can observe how for larger values, there is a minor performance drop, but generally RbM is relatively insensitive to the annealing schedule. We assume that large values of $\Delta$ potentially lead to the algorithm becoming stuck at poorer expert embeddings. Empirically, we observe that, even with $\Delta=0$, the expert embeddings converge sufficiently such that all representations are consistently routed to exactly the same experts in successive epochs.
Table 10: Routing-by-Memory with annealing.
Dataset Eval $\Delta=0.0$ $\Delta=0.05$ $\Delta=0.1$ $\Delta=0.5$ $\Delta=0.9$
OGB-ArXive ind 71.31$\pm$0.20 71.31$\pm$0.20 71.27$\pm$0.24 71.26$\pm$0.27 71.27$\pm$0.42
tran 72.48$\pm$0.13 72.48$\pm$0.13 72.46$\pm$0.22 72.44$\pm$0.20 72.34$\pm$0.25